Physics 115A xxxxx Fall 2013 Statistical Physics due 9:30 pm Tuesday, October 22

PROBLEM SET 3

Reading: Chapter 4 in Reif. Problems:

Reif 3.5 Reif 3.6 (Think about your answer to 3.5 to do this problem.) Reif 4.1 Reif 4.3 Reif 4.4 (Hint: Use the definition of entropy to calculate the entropy at T = ∞ and T = 0 and then use Eq. 4.5.1.)